OpenStudy (anonymous):

can sin^2A-cos^2A=-1?

OpenStudy (anonymous):

now it can't be we alredy know tht sin^2 + cos^2 =1 replc value to get answer

OpenStudy (anonymous):

lets rplc cos^2 cos^2 = 1 - sin^2 sin^2 - 1 + sin^2 2sin^2 - 1 which is not equal to -1

OpenStudy (anonymous):

\[\sin ^{2}\theta -\cos ^{2}\theta= -1\] \[\sin^{2}\theta=-(1-\cos ^{2}\theta)\] \[\sin ^{2}\theta=-\sin ^{2}\theta\] this implies that for this equation \[\sin \theta\] has fixed value 0